Caro Manufacturing has two production departments, Machining and Assembly, and two service departments, Maintenance and Cafeteria. Direct costs for each department and the proportion of service costs used by the various departments for the month of August follow:

Proportion of Services Used by
Department Direct Costs Maintenance Cafeteria Machining Assembly
Machining $ 140,000
Assembly 68,000
Maintenance 52,000 0.3 0.3 0.4
Cafeteria 33,000 0.8 0.1 0.1

Required:

Use the reciprocal method to allocate the service costs. (Matrix algebra is not required.) (Negative amounts should be indicated by a minus sign. Do not round intermediate calculations.)

Maintenance Cafeteria Machining Assembly
Service Department Costs
Maintenance
Cafeteria
Total

Thank You!
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Caro Manufacturing
Reciprocal Method Maintenance Cafeteria Machining Assembly Total
Maintenance                      0.30                        0.30                0.40                1.00
Cafeteria                             0.80                        0.10                0.10                1.00
Direct Costs                   52,000.00            33,000.00            140,000.00      68,000.00 293,000.00
Let Total cost of Maintenance is $ X and Total cost of Cafeteria is $ Y
So, X= 52000+ 0.80 of Y Equation 1
Y= 33000+ 0.30 of X Equation 2
Or, X= 52000+ 0.80Y
Y= 33000+ 0.30X
Substituting the value of X in equation 2:
Y = 33000 + 0.30(52,000 + 0.80Y)
Y = 33000 + 15,600+0.24Y
Y – 0.24Y = 33000 + 15,600
0.76Y= 48600
Y = 48600/0.76
Y = $ 63,947.37
Substituting the value of Y in equation 1:
So, X= 52000+ 0.80 of Y
So, X= 52000+ 0.80 of 63,947.37
So, X= 52000+ 51157.89
So, X= 103,157.89
Allocation Summary Maintenance Cafeteria Machining Assembly Total
Maintenance                      0.30                        0.30                0.40                1.00
Cafeteria                             0.80                        0.10                0.10                1.00
Direct Costs                   52,000.00            33,000.00            140,000.00      68,000.00 293,000.00
Allocation
Human Resources               (103,157.89)            30,947.37              30,947.37      41,263.16                    -  
General Factory                   51,157.89          (63,947.37)                6,394.74        6,394.74                    -  
Total cost after allocation                             0.00                    (0.00)           177,342.10 115,657.89 293,000.00